OpenStudy (anonymous):

What is the simplified form of √(8)(√(5) + 4)?

OpenStudy (lgbasallote):

\[\large \sqrt 8(\sqrt 5 + 4)\] is this the question?

OpenStudy (anonymous):

Yes, but it had parenthesis between the 5 as well. It might have been a typo from my teacher, though.

OpenStudy (lgbasallote):

that is equal to \[\sqrt 8 \times \sqrt 5 + \sqrt 8 \times 4\] if we simplify that we get \[\sqrt{40} + 4\sqrt 8\] simplify again \[\sqrt{4 \times 10} + 4\sqrt{4\times 2}\] another simplify... \[2\sqrt{10} + (4\times 2)\sqrt{2}\] last simplify... \[2\sqrt{10} + 8\sqrt 2\]

OpenStudy (anonymous):

Thank you so much! :D